# Step-Calibrated Diffusion for Biomedical Optical Image Restoration

### v1

@@ Appendix 0.B Details about SRH images
Appendix 0.C Methodology Additional Details
0.C.1 Step Calibrator Training and Calibration Distribution
We show the full algorithm of training the step calibrator with manually added Gaussian noise and augmentations in Algorithm 2, as well as one example of the augmentation in Figure 11.
@@ 0.C.1 Step Calibrator Training and Calibration Distribution
0.C.2 Diffusion Model Training
We show the full training algorithm in Algorithm 3 below. It is almost identical to the training algorithm from [15], except that we only sample up to .
Appendix 0.D Experiment details and additional results
0.D.1 Baseline and Ablation details
@@
CycleGAN [46]: We trained a CycleGAN model for image-to-image translation between unpaired low-quality and high-quality images. We used the 4.5K hand-picked high-quality and 3.8K hand-picked low-quality images for training the CyleGAN. Both the high-quality and low-quality image sets contained at least 10 different tumor types. The CycleGAN model performed better when trained on the manually selected images compared to training on the full 840K filtered high-quality images.
Synthetic noise[26, 30]: A U-Net model is trained by adding Gaussian noise to high-quality images to generate synthetic low-quality images. We train only with the 4.5K manually selected high-quality images because we found that the model performs better compared to training on the full 840K filtered high-quality images.
Conditional Diffusion [25, 36]: We train a conditional generative diffusion model, where we sample from a random Gaussian prior, and generate the restored image conditioned on the low-quality image. The conditioning method is concatenation (i.e. the input image and the conditioning image is concatenated as input to the UNet). We train the model with the 840K HQ training images and fine-tune it with the 4.5K hand-picked images. During training, we add a random amount of Gaussian noise to the high-quality images as conditioning images. We show the full training algorithm in Algorithm.
Regularized Reverse Diffusion (RRD) [5] is a recent method developed for restoring noisy medical images. It also uses part of the reverse diffusion process as restoration, but it employs a non-parametric noise estimator [3] to determine the step numbers, and uses a Fourier-feature-based regularization to reduce hallucination. Since we were unable to find any official code for the RRD paper and we deal with images in a vastly different domain from theirs (RRD was originally designed for denoising single-channel MRI images), we trained our own diffusion model following the same process as [15] and using the same cosine noise schedule as RSCD.
Deep Image Prior [38] is an unsupervised image restoration method that works by training a ResNet to reconstruct the noisy image but takes the generation of the model before its convergence as the denoised image. We take the output of step 1800 as our baseline, as images generated from this training step tend to have the best quality.
@@
No Dynamic Recalibration: RSCD without Dynamic Recalibration, i.e. Algorithm 1 without the blue parts. We take the initial step calibration and directly run that many steps of denoising to the end.
Nonparametric Noise Estimator: We replace our ResNet-based step calibrator with the nonparametric noise estimator [3] used in RRD [5]. We determine the step number as the closest cumulative noise to the predicted noise level from the nonparametric noise estimator.
@@
No Augmentation Step Calibrator: We use a step calibrator that is trained without augmentations (i.e. Algorithm 2 without the red parts).
0.D.2 Unpaired Images Evaluation additional qualitative results
@@
0.D.3 Near-registered dataset collection details
Imaging with low laser power is one of the known ways to produce noisy images. We generate a paired set of near-registered SRH images by imaging a surgical specimen immediately after the SRH imaging system is started (without laser warm-up and optimization), and then image the same specimen after 20 minutes, once the laser has been optimized and sufficiently warmed up. Next, image registration is required to spatially align the low-quality and high-quality images because the tissue specimen may be shifted due to gravity between the two image acquisitions. Paired images were registered by fitting an affine transformation using scale-invariant feature transform [28] and RANSAC [7].
@@ 0.D.4 Hyperparameters
Step Calibrator Training | Base Model | ResNet-50 | ||
@@
Learning Rate | 0.00001 | |||
Optimizer | Adam | |||
Epochs | 500 | |||
Batch Size | 150 | |||
Diffusion Model Architecture | Model | UNet | ||
In channels | 2 | |||
Init dim | 128 | |||
Out channels | 2 | |||
Dim mult | [1,2,2,4] | |||
Resnet block groups | 8 | |||
Convnext Mult | 2 | |||
Diffusion Model Training | Noise Schedule | Cosine (s=0.008) | ||
Total Steps (T) | 1000 | |||
200 | |||
Batch Size | 8 | |||
Learning Rate | 0.00001 | |||
Loss type | L1 | |||
@@
Optimizer | Adam | |||
Dynamic Recalibration | Recalibration interval (d) | 10 |
0.D.5 Computing Resource and Inference Efficiency
@@
0.D.6 Human Expert Preference Details
We show the human expert preference collection interface in Figure 14. The expert is given the original LQ image as well as 2 restorations (one from RSCD and one from a baseline, in random order; the expert is not told which restoration is baseline and which one is RSCD), and then the expert is asked to (1) select one restoration that the expert believes has better overall quality, and (2) determine which restoration has clinically significant hallucinations (can choose "neither" if neither restoration hallucinates). We perform a total of 1200 such pair-wise preference collections (100 LQ images x 6 baseline comparisons x 2 preferences from different experts per comparison), distributed among all participating experts.
Appendix 0.E Downstream task details
0.E.1 Deep learning-based tumor diagnosis
@@
A portion of our SRH images have ground truth labels on whether the tissue is normal or tumor. Therefore, we took the intersection of that labeled portion and the 12K unpaired low-quality images used for the experiment in section 4.2 as the dataset for deep learning-based tumor diagnosis. The dataset contains 8,372 low-quality images, each with its ground truth label of tumor/normal.
@@
We used the segmentation model from [20], the most influential and widely accepted study on deep learning-based brain tumor diagnosis through SRH images, as our classification model. The model classifies each SRH image into one of three classes: normal, tumor, and non-diagnostic. The model was trained to classify an SRH image as non-diagnostic if the model is unable to extract sufficient information from the image to make a diagnosis, and there are several possibilities for an image to be non-diagnostic, including: (1) the image scanning an irrelevant part of the slide, or (2) the image quality is too low.
@@
As shown in Table 7, the restored images have a significantly higher “correct" rate while having a much lower percentage of misclassified or non-diagnostic images. This means that RSCD can restore a significant portion of the images that were originally classified as non-diagnostic or wrongly classified due to poor quality, and make them high-quality enough to be correctly classified.
@@
0.E.2 Z-stack Image Restoration

### v2

# Supplemental Material Appendix A Additional Related Works A.1 Image Denoising and Restoration

@@ Appendix B Details about SRH images
[width=]figs/Fig_10.pdf
Appendix C Methodology Additional Details
C.1 Step Calibrator Training and Calibration Distribution
We show the full algorithm of training the step calibrator with manually added Gaussian noise and augmentations in Algorithm C.1, as well as one example of the augmentation in Figure 2.
[…]
@@ C.1 Step Calibrator Training and Calibration Distribution
[width=0.6]figs/aug.png
[width=0.6]figs/histogram.png
C.2 Diffusion Model Training
We show the full training algorithm in Algorithm C.2 below. It is almost identical to the training algorithm from [Ho2020-tm], except that we only sample up to .
[…]
{algorithmic}[1]
[…]
Appendix D Experiment details and additional results
D.1 Baseline and Ablation details
@@
CycleGAN [Zhu2017-wf]: We trained a CycleGAN model for image-to-image translation between unpaired low-quality and high-quality images. We used the 4.5K hand-picked high-quality and 3.8K hand-picked low-quality images for training the CyleGAN. Both the high-quality and low-quality image sets contained at least 10 different tumor types. The CycleGAN model performed better when trained on the manually selected images compared to training on the full 840K filtered high-quality images.
Synthetic noise[Lehtinen2018-zd, Manifold2019-ft]: A U-Net model is trained by adding Gaussian noise to high-quality images to generate synthetic low-quality images. We train only with the 4.5K manually selected high-quality images because we found that the model performs better compared to training on the full 840K filtered high-quality images.
Conditional Diffusion [Kawar2022-vc, Saharia2022-co]: We train a conditional generative diffusion model, where we sample from a random Gaussian prior, and generate the restored image conditioned on the low-quality image. The conditioning method is concatenation (i.e. the input image and the conditioning image is concatenated as input to the UNet). We train the model with the 840K HQ training images and fine-tune it with the 4.5K hand-picked images. During training, we add a random amount of Gaussian noise to the high-quality images as conditioning images. We show the full training algorithm in Algorithm.
Conditional Diffusion: Training
{algorithmic}[1]
[…]
Regularized Reverse Diffusion (RRD) [Chung2023-qt] is a recent method developed for restoring noisy medical images. It also uses part of the reverse diffusion process as restoration, but it employs a non-parametric noise estimator [Chen2015-ij] to determine the step numbers, and uses a Fourier-feature-based regularization to reduce hallucination. Since we were unable to find any official code for the RRD paper and we deal with images in a vastly different domain from theirs (RRD was originally designed for denoising single-channel MRI images), we trained our own diffusion model following the same process as [Ho2020-tm] and using the same cosine noise schedule as RSCD.
Deep Image Prior [Ulyanov2016-kz] is an unsupervised image restoration method that works by training a ResNet to reconstruct the noisy image but takes the generation of the model before its convergence as the denoised image. We take the output of step 1800 as our baseline, as images generated from this training step tend to have the best quality.
@@
No Dynamic Recalibration: RSCD without Dynamic Recalibration, i.e. Algorithm LABEL:alg:dr without the blue parts. We take the initial step calibration and directly run that many steps of denoising to the end.
Nonparametric Noise Estimator: We replace our ResNet-based step calibrator with the nonparametric noise estimator [Chen2015-ij] used in RRD [Chung2023-qt]. We determine the step number as the closest cumulative noise to the predicted noise level from the nonparametric noise estimator.
@@
No Augmentation Step Calibrator: We use a step calibrator that is trained without augmentations (i.e. Algorithm C.1 without the red parts).
D.2 Unpaired Images Evaluation additional qualitative results
@@
[width=]figs/supp.unpaired.pdf
D.3 Near-registered dataset collection details
Imaging with low laser power is one of the known ways to produce noisy images. We generate a paired set of near-registered SRH images by imaging a surgical specimen immediately after the SRH imaging system is started (without laser warm-up and optimization), and then image the same specimen after 20 minutes, once the laser has been optimized and sufficiently warmed up. Next, image registration is required to spatially align the low-quality and high-quality images because the tissue specimen may be shifted due to gravity between the two image acquisitions. Paired images were registered by fitting an affine transformation using scale-invariant feature transform [Lowe2004-gq] and RANSAC [derpanis2010overview].
@@ D.4 Hyperparameters
[…]
Base Model | ResNet-50 | |||
@@
| Learning Rate | 0.00001 | ||||
| Optimizer | Adam | ||||
| Epochs | 500 | ||||
| Batch Size | 150 | ||||
[…]
Model | UNet | |||
| In channels | 2 | ||||
| Init dim | 128 | ||||
| Out channels | 2 | ||||
| Dim mult | [1,2,2,4] | ||||
| Resnet block groups | 8 | ||||
| Convnext Mult | 2 | ||||
[…]
Noise Schedule | Cosine (s=0.008) | |||
| Total Steps (T) | 1000 | ||||
200 | ||||
| Batch Size | 8 | ||||
| Learning Rate | 0.00001 | ||||
| Loss type | L1 | ||||
@@
| Optimizer | Adam | ||||
[…]
D.5 Computing Resource and Inference Efficiency
@@
D.6 Human Expert Preference Details
We show the human expert preference collection interface in Figure 5. The expert is given the original LQ image as well as 2 restorations (one from RSCD and one from a baseline, in random order; the expert is not told which restoration is baseline and which one is RSCD), and then the expert is asked to (1) select one restoration that the expert believes has better overall quality, and (2) determine which restoration has clinically significant hallucinations (can choose ”neither” if neither restoration hallucinates). We perform a total of 1200 such pair-wise preference collections (100 LQ images x 6 baseline comparisons x 2 preferences from different experts per comparison), distributed among all participating experts.
[width=]figs/interface.pdf
Appendix E Downstream task details
E.1 Deep learning-based tumor diagnosis
@@
A portion of our SRH images have ground truth labels on whether the tissue is normal or tumor. Therefore, we took the intersection of that labeled portion and the 12K unpaired low-quality images used for the experiment in section LABEL:sec:unpaired as the dataset for deep learning-based tumor diagnosis. The dataset contains 8,372 low-quality images, each with its ground truth label of tumor/normal.
@@
We used the segmentation model from [Hollon2020-ez], the most influential and widely accepted study on deep learning-based brain tumor diagnosis through SRH images, as our classification model. The model classifies each SRH image into one of three classes: normal, tumor, and non-diagnostic. The model was trained to classify an SRH image as non-diagnostic if the model is unable to extract sufficient information from the image to make a diagnosis, and there are several possibilities for an image to be non-diagnostic, including: (1) the image scanning an irrelevant part of the slide, or (2) the image quality is too low.
@@
As shown in Table LABEL:fig:classifier, the restored images have a significantly higher “correct” rate while having a much lower percentage of misclassified or non-diagnostic images. This means that RSCD can restore a significant portion of the images that were originally classified as non-diagnostic or wrongly classified due to poor quality, and make them high-quality enough to be correctly classified.
@@
[width=]figs/Fig_15.pdf
E.2 Z-stack Image Restoration